OpenStudy (anonymous):

The sum of three consecutive even numbers is 44 more than the average of these numbers. The product of the smallest and the largest numbers is equal to

OpenStudy (anonymous):

yes i got 437.5

OpenStudy (anonymous):

x+x+2+x+4=44+x x=19, largest =23

OpenStudy (anonymous):

x = 1st integer x + 2 = 2nd integer x + 4 = 3rd integer x + x + 2 + x + 4 = 44 + (x + x + 2+ x + 4)/3 3x + 6 = 44 + (3x + 6)/3--> (3x + 6/3 = x + 2) 3x + 6 = 44 + x + 2 3x + 6 = 46 + x 3x - x = 46 - 6 2x = 40 x = 20 x + 2 = 20 + 2 = 22 x + 4 = 20 + 4 = 24 check... add 20 + 22 + 24 = 66. ----> 66/3 = 22 (average) The sum is 44 more then the average. The product of the smallest and the largest number is : 20 x 24 = 480

OpenStudy (anonymous):

Do you understand what I just did ?

OpenStudy (anonymous):

yes i got it .thanks

OpenStudy (anonymous):

your welcome...glad to help :)

OpenStudy (anonymous):

what do you mean by qualifications ?

OpenStudy (anonymous):

What will be the difference between the simple and the compound interest at the rate of 15% per annum on a sum of Rs 12000 after 3 years?

OpenStudy (anonymous):

See the attached for a slightly different solution approach.

OpenStudy (anonymous):

F o u r p i p e s fill a t a n k in 1 5, 2 0 , 30 a n d 40 h o u rs respectively. T h e first pipe wa s o p e n e d at 6 am, the second at 8 am, the third at 10 am and the fourth at 12 noon. When will the tank get filled completely?

OpenStudy (anonymous):

In the first 6 hours, 0600 to 1200, the first, second and third pipe fill 2/3 of the tank together.\[\frac{6}{15}+\frac{4}{20}+\frac{2}{30}=\frac{2}{3} \]Solve the following for t, the time in hours for the remaining 1/3 of the tank to be fillled with all 4 pipes open.\[\frac{1}{15}+\frac{1}{20}+\frac{1}{30}+\frac{1}{40}=\frac{1}{3 t} \] The RHS is "one third divided by t".\[t=\frac{40}{21}=\text{ 1 hr 54.3 minutes after noon} \]